2. Calculate the molar mass of the solute in a solution of 2.47 g of a mo electrolyte in 100 g of acetic acid (CH3COO). The solution freezes at 1.3 °C below the normal freezing point of pure acetic acid. (Kf for CH3COOH=3.9°C/m)


1
Expert's answer
2020-10-21T13:20:08-0400

2.47 g of a mo electrolyte


100 g of acetic acid (CH3COO).


The solution freezes at 1.3 °C

(Kf for CH3COOH=3.9°C/m)


Therefore :


2.47/ 100 =0.0247moles


1 mole. = 3.9 OC


? = 1.3 OC

= 1.3/3.9 = 0.3moles


0.0247moles = 2.47g


0.3 moles = ((0.3)(2.47))/0.0247


= 30g